# Towards Optimal Environmental Policies: Policy Learning under Arbitrary Bipartite Network Interference

### v2

@@ E.3 Count Analysis
Figure E.3: Grid with the reduction of IHD Hospitalizations, varying the spending from 10%-90% of budget
Appendix F Model Generalizations and Extensions
Multiple Timepoints
@@ Appendix F Model Generalizations and Extensions
If we have $n$ samples over $t\in[T]$ timepoints, we may follow the same approach as in Schulte et al. (2014). For illustration purposes, we once again assume the outcome model at each timepoint factorizes by intervention unit as in Equation 5:
@@
$=f_{0t}(\mathbf{X}_{it}^{out},\boldsymbol{\alpha}_{0t})+\bar{a}_{it}f_{At}(\mathbf{X}_{it}^{out},\boldsymbol{\beta}_{0t})+\epsilon_{it}$ (F.1)
Above, each quantity is the analog as described in Equation 5, but indexed a second time for timepoint $t\in[T]$.
Q-Learning
@@
$\frac{1}{n}\sum_{i=1}^{n}\phi_{t}(Y_{it};\boldsymbol{\theta})=0$ (F.2)
@@
Specifically, at any timepoint $t$, we obtain $\hat{\pi}^{*}_{t}(\widehat{\texttt{TotalEffect}_{jt}})$. Thus at time $t-1$, we have the optimal policy value assuming policymakers make the optimal policy choice at time $t$. With backwards induction, model-correctness, and straightforward extensions of the assumptions mentioned throughout Section 2, this strategy can be shown to be valid.
A-Learning
The result follows similarly to above. Extending the estimating equations in question Equation 11, Equation 12, Equation 10 to $T$ timepoints and solving for these backwards in time yields a multi-time point extension to A Learning.
Nonparametric Models
The approaches above can be extended to more flexible models assuming nuisance parameters are estimated at $\sqrt{n}$ consistent rate. Valid inference would follow from Section 2.3.1 and Theorem 1 since higher order terms would decay to 0 fast enough. In particular, one may employ nonparametric series estimators for functions that satisfy relevant smoothness conditions (Giné & Nickl, 2015).
Mean Outcome Form
The proposed framework permits use of more general mean outcome models than that considered in Equation 5. Equation 5 enables us to obtain closed-form or efficient solutions to Equation 4. In general, we may require higher computational burden depending on the assumed outcome form. Interpretability also goes down since we would not be able to directly interpret TotalEffect in terms of the outcome model as in Equation 6.

### v3

@@ 9 Supplementary Material
Figure Appendix N.3: Count Results. Grid with the reduction of IHD Hospitalizations, varying the spending from 10%-90% of budget.
Appendix N.4 Optimal policy building upon the factual 2005 scrubber landscape
In this section, we consider the alternative to analysis presented in Section 4, examining policies starting from the factual scrubber distribution, only allowing for addition of new scrubbers. We refer to this as the augmentation policy. We do not consider the possibility of removing existing scrubber installations because (1) we do not have data on the cost and effects of doing so, and (2) policymakers may not consider such an action since it may be unethical given the established links between air pollution and adverse public health outcomes. Because the factual scrubber policy costs approximately 10% of the universal scrubber installation cost, we consider the optimal cost-constrained policy under 20%, 30% …90% of the universal cost in Fig. Appendix N.4. The factual policy is listed in the 10% slot, and the 100% treated policy is identical to the 100% policy in our main analysis.
Refer to caption Refer to caption Refer to caption
(a) 10% Total Cost (b) 20% Total Cost (c) 30% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 40% Total Cost (b) 50% Total Cost (c) 60% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 70% Total Cost (b) 80% Total Cost (c) 90% Total Cost
Figure Appendix N.4: Grid with the reduction of IHD Hospitalizations/10,000 Person Years, varying the spending from 10%-90% of budget, when considering the augmentation policy.
For further comparison, we contrast the clean-slate analysis and augmentation policy in Table Appendix N.2. We see that the augmentation policy is inefficient along the range of cost constraints considered, with notable gains experienced using the clean slate analysis under lower cost constraints. As we increase the cost to around 40-100%, the rate reductions are very similar.
%$C$ Clean Slate Augmentation
10% 20.9 6.1
20% 29.0 23.7
30% 34.9 31.8
40% 39.0 37.1
50% 42.2 40.9
60% 44.4 43.6
70% 45.9 45.3
80% 46.8 46.3
90% 47.4 46.9
Table Appendix N.2: Rate reductions by cost constraint (% of universal scrubber cost $C$) for the clean-slate analysis and the augmentation policy.
Appendix N.5 Optimal Policy vs. Factual Policy
In this subsection, we compare the factual 2005 scrubber installation scheme and the optimal “clean slate” policy under the same cost constraint ($\sim 10\%$ of the universal cost). As shown in Table Appendix N.2 and Fig. Appendix N.5, the clean-slate analysis yields notable improvement in IHD hospitalization reductiosn compared to the factual policy. We can see on the map in Fig. Appendix N.5, the factual policy treats 73 plants, compared to the optimal clean-slate policy, which treats 117 plants in similar regions and yields higher IHD hospitalization rate reductions. This indicates lower efficiency in the factual policy, since it chooses more expensive and less health benefiting power plants. In fact, there are 27 power plants overlapping between the optimal policy and factual policy, so the factual policy did not choose many that the optimal policy would have chosen under this cost constraint.
Refer to caption Refer to caption
(a) Empirical Policy (b) Optimal Policy
Figure Appendix N.5: Empirical and Optimal Policy Comparison, under equivalent budget constraints.
Appendix N.6 Sensitivity Analyses on Plants without Scrubbers newly installed in 2005.
Of the 74 power plants in our dataset with scrubber installations by 2005, only 4 (1%) of these power plants had scrubbers newly installed in year 2005. This raises the potential concern for temporality of outcomes occurring prior to treatment since outcomes were measured in 2005. In this section, we carry out sensitivity analyses excluding these 4 power plants to test how sensitive our results may be to potential violation of this assumption. Specifically, we analyze how our $\texttt{TotalEffect}_{j}$ estimation differs.
Rerunning our analysis without these plants, we obtain nearly identical rate and count reductions (identical up to 0.1 for rates, and identical up to 1 for counts) – Fig. Appendix N.6 supports this claim showing nearly identical histogram plots of $\widehat{\texttt{TotalEffect}}_{j}$.
Based on this analysis, we have reasonable confidence that the temporal, causal order is respected since the majority of power plants have scrubbers installed prior to year 2005 (when we measure outcomes) and the results are not sensitive to exclusion of these 4 power plants.
Refer to caption Refer to caption
(a) Include 4 plants (current analysis) (b) Exclude 4 plants (sensitivity analysis)
Figure Appendix N.6: Comparison of $\widehat{\texttt{TotalEffect}}_{j}$ when we include the 4 plants that have scrubbers installed in the year 2005 (left), and exclude the 4 plants (right).
Appendix O Implementing the Policy on Future Units
Suppose we were interested in recommending scrubber installations for a new set of power plants. In order to carry out this analysis, we will need an updated, or even possibly new, bipartite adjacency matrix $\mathbf{H}^{\prime}$ (HyADS matrix) that holds information on the dependencies between the new units $(\mathbf{X}^{int}_{j})^{\prime},(\mathbf{X}^{out}_{i})^{\prime}$ yielding sample sizes $J^{\prime},n^{\prime}$. Then we can take the following approaches under no cost-constraints and cost-constraints.
Appendix O.1 Policies under No Cost Constraints.
Recall the policy for power plant $j$ under no cost constraints is given by
$\hat{\pi}_{j}^{*}(\texttt{TotalEffect}(\hat{\boldsymbol{\beta}})_{j})=\textbf{1}\{\texttt{TotalEffect}(\hat{\boldsymbol{\beta}})_{j}<0\}$
For a future power plant $j^{\prime}$, we can estimate the policy in the same manner:
$\hat{\pi}_{j^{\prime}}^{*}(\texttt{TotalEffect}(\hat{\boldsymbol{\beta}})_{j^{\prime}})=\textbf{1}\{\texttt{TotalEffect}(\hat{\boldsymbol{\beta}})_{j^{\prime}}<0\}$ (Appendix O.1)
[…]
Appendix O.2 Policies under Cost Constraints.
Recall that under cost-constraints, we solve Eq. 4. This is equivalent to choosing to treat power plants with largest $\widehat{\texttt{BenefitCost}}_{j}$ until our cost-constraint $C$ is hit. Denote the rank of power plant $j$’s $\widehat{\texttt{BenefitCost}}_{j}$ by $(j)$. Denote the last power plant that is able to be treated under cost constraint $C$ by $j^{*}_{C}$. Then the power plant that has the next highest $\widehat{\texttt{BenefitCost}}$ (power plant $(j^{*}_{C})+1$) is treated with probability proportional to the remaining budget. In other words, the policy is given by:
$\hat{\pi}_{j}^{*}(\texttt{BenefitCost}(\hat{\boldsymbol{\beta}})_{j})=\begin{cases}1,&\text{if }(j)\leq(j^{*}_{C})\\ \frac{C-\sum_{l:l\leq(j^{*}_{C})}c_{l}}{c_{j}},&\text{if }(j)=(j^{*}_{C})+1\\ 0,&\text{Otherwise.}\end{cases}$
Then, we may consider a similar adaptation to policy making on a future power plant $j^{\prime}$, recalculating the last power plant that is able to be treated under cost-constraint $C$ accounting for $j^{\prime}$. Denote this power plant by $j^{**}_{C}$.
$\hat{\pi}_{j^{\prime}}^{*}(\texttt{BenefitCost}(\hat{\boldsymbol{\beta}})_{j^{\prime}})=\begin{cases}1,&\text{if }(j)\leq(j^{**}_{C})\\ \frac{C-\sum_{l:l\leq(j^{**}_{C})}c_{l}}{c_{j}},&\text{if }(j)=(j^{**}_{C})+1\\ 0,&\text{Otherwise.}\end{cases}$
Appendix P Model Generalizations and Extensions
Appendix P.1 Multiple Timepoints
@@
If we have $n$ samples over $t\in[T]$ timepoints, we may follow the same approach as in Schulte et al. [2014]. For illustration purposes, we once again assume the outcome model at each timepoint factorizes by intervention unit as in Eq. 5:
@@
$=f_{0t}(\mathbf{X}_{it}^{out},\boldsymbol{\alpha}_{0t})+\bar{a}_{it}f_{At}(\mathbf{X}_{it}^{out},\boldsymbol{\beta}_{0t})+\epsilon_{it}$ (Appendix P.1)
Above, each quantity is the analog as described in Eq. 5, but indexed a second time for timepoint $t\in[T]$.
Appendix P.2 Q-Learning
@@
$\frac{1}{n}\sum_{i=1}^{n}\phi_{t}(Y_{it};\boldsymbol{\theta})=0$ (Appendix P.2)
@@
[…]
Appendix P.3 A-Learning
The result follows similarly to above. Extending the estimating equations in question Eq. 10, Eq. 11, Eq. 9 to $T$ timepoints and solving for these backwards in time yields a multi-time point extension to A Learning.
Appendix P.4 Nonparametric Models
The approaches above can be extended to more flexible models assuming nuisance parameters are estimated at $\sqrt{n}$ consistent rate. Valid inference would follow from Section 2.3.2 and Theorem 3 since higher order terms would decay to 0 fast enough. In particular, one may employ nonparametric series estimators for functions that satisfy smoothness conditions [Giné and Nickl, 2015].
Appendix P.5 Mean Outcome Form
The proposed framework permits use of more general mean outcome models than that considered in Eq. 5. Eq. 5 enables us to obtain closed-form or efficient solutions to Eq. 4. In general, we may require higher computational burden depending on the assumed outcome form. Interpretability also goes down since we would not be able to directly interpret TotalEffect in terms of the outcome model as in Eq. 6.